### v1

# Mitigating the Impact of Outlier Channels for Language Model Quantization with Activation Regularization

Abstract
We consider the problem of accurate quantization for language models, where both the weights and activations are quantized to 4 bits per parameter with uniform quantization, the lowest bitwidth format natively supported by existing GPU hardware. In this context, the key challenge is activation quantization: it is known that language models contain outlier channels whose values on average are orders of magnitude higher than than other channels, which prevents accurate low-bitwidth quantization with known techniques. We systematically study this phenomena and find that these outlier channels emerge early in training, and that they occur more frequently in layers with residual streams. We then propose a simple strategy which regularizes a layer’s inputs via quantization-aware training (QAT) and its outputs via activation kurtosis regularization. We show that regularizing both the inputs and outputs is crucial for preventing a model’s “migrating” the difficulty in input quantization to the weights, which makes post-training quantization (PTQ) of weights more difficult. When combined with weight PTQ, we show that our approach can obtain a W4A4 model with integer quantization that performs competitively to the standard-precision W16A16 baseline. [^1] Code is available at https://github.com/aninrusimha/qat-pretrain
Aniruddha Nrusimha1Mayank Mishra2Naigang Wang3
Dan Alistarh4,5Rameswar Panda2Yoon Kim1
1Massachusetts Institute of Technology 2MIT-IBM Watson AI Lab
3IBM Research 4IST Austria 5NeuralMagic
anin@mit.edu
1 Introduction
Large language models (LLM) have been shown to contain outlier channels, i.e., feature dimensions whose values are orders of magnitude higher than the others. These outlier channels are known to be crucial for strong model performance (Kovaleva et al., 2021; Puccetti et al., 2022), but pose significant challenges from a model compression perspective, for instance via post-training quantization (PTQ) (Dettmers et al., 2022; Xiao et al., 2023; Wei et al., 2022). Concretely, to enable the use of low-bitwidth integer matrix multiplications—which can lead to significant speed-ups—both the activations and the weights need to be quantized. However the presence of high outlier values in the model activations results in high quantization errors, and thus overall poor PTQ accuracy (see, e.g., Xiao et al. (2023)).
To mitigate the effect of outlier channels for activation quantization at the per-tensor level, existing works have explored various approaches, including keeping some of the computations in higher precision (Dettmers et al., 2022; Ashkboos et al., 2023; Zhao et al., 2023), or “migrating” the difficulty of quantizing outlier channels to other parts of the model (Xiao et al., 2023; Wei et al., 2023; Liu et al., 2023). While the above strategies have been effective for achieving INT8 activation quantization, INT4 quantization with PTQ methods remains an open challenge, with current methods still facing nontrivial degradations in perplexity (Wu et al., 2023; Shao et al., 2023; Yuan et al., 2023).
Figure 1: (Top) Average of the absolute activation values of a KV projection layer for a 1B language model trained with (a) standard training, (b) QAT with learned clipping values in the input layer, and (c) QAT on the inputs and kurtosis regularization on the layer’s outputs. For the QAT runs, we show the learned clip value as a green 2d manifold. (Bottom) Parameter values of individual weights in the KV projection of the same layer corresponding to each model after training. QAT-only training results in the model’s weights’ becoming harder to quantize, whereas kurtosis regularization mitigates this.
In this work, we perform an empirical study of outlier channel phenomena from a pretraining perspective. We find that dimensions with outlier channels emerge relatively early in training (see fig. 1(a), top), suggesting that their mitigation requires early intervention. These outlier channels are particularly prevalent in the output projection layer of the first layer, as well as the query-key-value projection layers of the other layers. Next, we explore a simple strategy that regularizes a layer’s input and output. On the input side, we show that a quantization-aware training (QAT) approach which learns the clipping values for each activation layer (Choi et al., 2018; Bhalgat et al., 2020) is effective at controlling the number of outlier channels, in addition to mitigating the effect of outliers through clipping (see fig. 1(b), top). However, while this approach can train a W16A4 model that has similar perplexity to a W16A16 model, post-training weight quantization to W4A4 results in nontrivial perplexity degradations, due to the model’s weights now becoming more difficult to quantize (see fig. 1(b), bottom). We thus additionally regularize the kurtosis of a layer’s output, which discourages the creation of outliers wholesale. Specifically, this discourages the layer’s weights having pathologically large rows (fig. 1(c), bottom).
Putting all these elements together, we show that we can train a language model at moderate scale (1 billion parameter models trained on 20 billion tokens) whose W4A4 perplexity is competitive to the standard-precision W16A16 baseline.
2 Background and Related Work
2.1 Uniform Quantization & Quantized Matmuls
We focus on uniform quantization, where the quantized values are evenly spaced between an interval range. Formally, for a given matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ that we wish to quantize to $b$ bits, let $c^{-}$ and $c^{+}$ be the pre-defined (or learned) clipping values. The quantization function $Q:\mathbb{R}^{n\times m}\to\mathbb{Z}^{n\times m}$ is then given by,
$Q({\mathbf{A}})=\operatorname{round}(s\times\operatorname{clamp}(\mathbf{A},c^{-},c^{+})+z),$
where $s=\frac{2^{b}-1}{c^{+}-c^{-}}$ is the scale factor and $z=\operatorname{round}(s\times c^{-})$ is the (optional) zero-point offset. This function, which can be generalized to different granularities of $\mathbf{A}$ (e.g., rows, columns or subgroups) transforms the entries of $\mathbf{A}$ into integers between $[0,2^{b}-1]$.
The quantized matrix $\mathbf{Q}_{\mathbf{A}}=Q(\mathbf{A})$ can be utilized in two different ways. First, the value can be dequantized to its original precision via $\widehat{\mathbf{A}}=\frac{1}{s}(\mathbf{Q}_{\mathbf{A}}-z)$ before multiplication. This method is typically used by pure weight quantization schemes, which multiply in the precision the model was trained in. Weight-only quantization can reduce a model’s memory footprint, and insofar as LLM inference is often memory bound, it can also enable faster inference by reducing the amount of time spent on memory operations during the forward pass (Lin et al., 2023; Frantar & Alistarh, 2024). However, the fact that the actual matmul is done in high precision is a fundamental limitation of weight-only quantization.
Second, the quantized values can be directly used for the matrix multiplication. Let $\mathbf{Q}_{\mathbf{U}}=Q(\mathbf{U})$, $\mathbf{Q}_{\mathbf{V}}=Q(\mathbf{V})$ be the quantized versions of $\mathbf{U}\in\mathbb{R}^{n\times k}$, $\mathbf{V}\in\mathbb{R}^{k\times m}$ with the respective scaling factors $s_{\mathbf{U}},s_{\mathbf{V}}$ and offsets $z_{\mathbf{U}},z_{\mathbf{V}}$. We can approximate $\mathbf{U}\mathbf{V}$ with
$\mathbf{U}\mathbf{V}$ $\approx\widehat{\mathbf{U}}\widehat{\mathbf{V}}=\frac{1}{s_{\mathbf{U}}s_{\mathbf{V}}}\times(\mathbf{Q}_{\mathbf{U}}-z_{\mathbf{U}})(\mathbf{Q}_{\mathbf{V}}-z_{\mathbf{V}}),$
where we can make use of low-precision matmuls for $(\mathbf{Q}_{\mathbf{U}}-z_{\mathbf{U}})(\mathbf{Q}_{\mathbf{V}}-z_{\mathbf{V}})$. In cases where the rows of $\mathbf{U}$ and columns of $\mathbf{V}$ are quantized separately with the corresponding scaling vectors $\mathbf{s}_{\mathbf{U}}\in\mathbb{R}^{n},\mathbf{s}_{\mathbf{V}}\in\mathbb{R}^{m}$ and offset vectors $\mathbf{z}_{\mathbf{U}}\in\mathbb{Z}^{n},\mathbf{z}_{\mathbf{V}}\in\mathbb{Z}^{m}$, we can still make use of integer matmuls since $\widehat{\mathbf{U}}\widehat{\mathbf{V}}$ is given by
$\operatorname{diag}(\mathbf{s}_{\mathbf{U}})^{-1}(\mathbf{Q}_{\mathbf{U}}-\mathbf{z}_{\mathbf{U}}\otimes\mathbf{1}_{k})(\mathbf{Q}_{\mathbf{V}}-\mathbf{1}_{k}\otimes\mathbf{z}_{\mathbf{V}})\operatorname{diag}(\mathbf{s}_{\mathbf{V}})^{-1}$
where $\mathbf{1}_{k}\in\mathbb{Z}^{k}$ is a vector of 1s and $\otimes$ is the outer product. [^2] If the offset vectors are not integers we can expand the expression and still use integer matmuls for $\mathbf{Q}_{\mathbf{U}}\mathbf{Q}_{\mathbf{V}}$. For the cross terms we can use the identity $(\mathbf{z}_{\mathbf{u}}\otimes\mathbf{1}_{k})\mathbf{Q}_{\mathbf{V}}=\mathbf{z}_{\mathbf{u}}\otimes(\mathbf{1}_{k}^{\top}\mathbf{Q}_{\mathbf{V}})$, and thus we can still make use of integer matmuls for most of the FLOPs. Note, however, lower-precision matmuls cannot straightfowardly be used if the $\mathbf{U}$ is quantized at the column level.
This second strategy which makes use of lower-precision matmuls can significantly improve inference latency and energy efficiency on supported hardware. For example, INT4 tensor core matmuls can be up to four times faster than FP16 tensor core matmuls on the NVIDIA Ampere architecture, [^3] https://developer.nvidia.com/blog/nvidia-ampere-architecture-in-depth/ while from a hardware-efficiency perspective, dedicated hardware for integer operations require much less area and energy usage than their floating-point counterparts (Jouppi et al., 2021; van Baalen et al., 2023).
2.2 Challenges in LLM Quantization
In LLMs, the majority of FLOPs are spent on dense matmuls of the form $\mathbf{X}\mathbf{W}$ where $\mathbf{X}\in\mathbb{R}^{L\times d_{in}}$ are the input activations (for $L$ input tokens) and $\mathbf{W}\in\mathbb{R}^{d_{in}\times d_{out}}$ are the model weights. For the Transformer architecture in particular this corresponds to the key, query, value projection layers, as well as the FFN layers. Given the sheer number of FLOPs in LLMs, inference efficiency can be improved significantly through lower-precision matmuls.
While there has been much work on post-training weight-only quantization for pretrained LLMs (Frantar et al., 2022; Dettmers & Zettlemoyer, 2023; Lin et al., 2023; Kim et al., 2023; Dettmers et al., 2023; Chee et al., 2023; Lee et al., 2023; Egiazarian et al., 2024, inter alia), PTQ for activations remains difficult due to the presence of outlier channels in LLMs trained with standard precision (Dettmers et al., 2022; Xiao et al., 2023). Informally, outlier channels are a set of input channels (i.e., columns of $\mathbf{X}$) whose values are many orders of magnitudes higher than the others, and have been shown to be crucial for performance (Kovaleva et al., 2021). If one were just interested in quantizing $\mathbf{X}$ independently, outlier channels could be managed by quantizing each column of $\mathbf{X}$ separately such that the scaling factor associated with an outlier channel is commensurate. However, as outlined in the previous section this would not enable the use of lower-precision matmuls, which requires $\mathbf{X}$ to be quantized by (at most) rows; unfortunately row-level (i.e., per-token) quantization results in significant performance degradations (Xiao et al., 2023).
2.3 Quantization-Aware Training
Figure 2: Frequency of outlier channels over the course of training. (Left) Proportion of outlier channels by layer depth. Layer 1 has highest occurrence of outlier channels. (Middle) In layer 1 inputs to the attention projection layer have the most outlier channels. (Right) This is generally not the case for the other layers, where the input to the QKV project layer has the most outlier channels.
Quantization-aware training (QAT) describes a class of techniques which aims to enable better quantization by simulating quantization during training (Zhou et al., 2016; Jacob et al., 2018; Zhang et al., 2018; Jung et al., 2019; Jain et al., 2020, inter alia). While there are many methods for QAT, we use a simple modified version of PACT (Choi et al., 2018) and LSQ (Bhalgat et al., 2020), which learn the clip values $c^{-}$ and $c^{+}$ for the activations. This approach uses the learned clip values to perform quantization during the forward pass, and uses the straight-through estimator for the gradients with respect to the clip values. While QAT has been studied extensively in the context of (typically smaller) vision models, QAT for pretraining language models with more than a billion parameters remains less explored.
Figure 3: Trajectory of a channel’s activations across 50B tokens of training. We show each channel’s absolute activation value averaged across 500K tokens.
3 Motivating Study: Outlier Channels in Language Models
We first conduct a preliminary analysis to study the emergence of outlier channels during pretraining, with both our own and open-source models. For our own pretrained models, we use the standard “pre-LayerNorm” Transformer architecture (Xiong et al., 2020), where given layer $l$’s input $\mathbf{X}^{(l)}\in\mathbb{R}^{L\times d}$ we obtain the next layer $\mathbf{X}^{(l+1)}$ via,
[…]
Here $\mathbf{W}_{QKV}\in\mathbb{R}^{d\times 3d},\mathbf{W}_{O}\in\mathbb{R}^{d\times d},\mathbf{W}_{1}\in\mathbb{R}^{d\times 4d},\mathbf{W}_{2}\in\mathbb{R}^{4d\times d}$ are learnable matrices, and the bias vectors are omitted for brevity. Our study focuses on the following activations that have been previously found to contain outlier channels: QKV Input ($\mathbf{Y}_{1}$), Attn Proj Input ($\mathbf{Y}_{2}$), MLP Input ($\mathbf{Y}_{3}$), MLP Proj Input ($\mathbf{Y}_{4}$). We train 1 billion parameter (24-layer model with 1920 dimensions) on 50 billion tokens from the SlimPajama dataset (Soboleva et al., 2023). We periodically collect activation statistics for all layers by running model checkpoints on (the same) 500K tokens from the C4 dataset.
First, we attempt to measure the prevalence of outlier channels aggregated by layer type and depth. For the purposes of this analysis, we name a channel an outlier if the average absolute value of the channel is over six times the average absolute value of all the input activations. This definition of an outlier channel is somewhat arbitrary, but similar definitions in the literature based on the other metrics (Kovaleva et al., 2021) generate similar results; we use this definition as opposed to definitions on the absolute values (Dettmers et al., 2022) to enable comparison across different layers. The results of this analysis are in fig. 2. Our results generally follow what has been established in the literature: while outliers are distributed throughout depth, the layers which tend to have the most outlier channels in their input are those whose inputs are the residual stream of the network. Interestingly, we find that outlier channels emerge early in training, and rapidly become numerous. The proportion of outlier channels within a layer then decreases gradually and eventually plateaus.
Figure 4: The distribution of activations over of a non-outlier channel (left) and two outlier channels (middle, right) over training.
Figure 5: Activation development in two open-source models: Pythia 6.9b (Biderman et al., 2023) and OLMo 7B (Groeneveld et al., 2024). We show activiations for a layer that reads from the residual stream (QKV Input) and one that does not (MLP Proj Input). Note that the OLMo data includes a step-0 checkpoint (i.e., at initialization).
We next perform a more granular analysis, where we analyze the average absolute value of channels over the training of a 1B model with 50B tokens. This is shown in fig. 3. Within channels, we observe that the development of outliers occurs early on during training. In most cases outliers primarily occur in layers that take as input the residual stream, although there is still significant variation in the average magnitude of channels in the input to other layers. We take a closer examination of the development of some the largest individual outlier channels for a particular layer in fig. 4. Channel 600, which is not an outlier channel, has channel values that are distributed roughly as a Gaussian with a mean of zero. The outlier channels, in comparison, have mean values that are significantly different from zero. This initial examination suggests that outlier channels are not scaled differently than non outlier channels, but have a shifted distribution. This potentially indicates why scaling and shifting methods, like OmniQuant (Shao et al., 2023), outperform scaling-only methods such as SmoothQuant (Xiao et al., 2023).
Open-source Models.
To validate the generality of our observations, we perform our analysis on two publicly available 7B models with public checkpoints, Pythia (Biderman et al., 2023) and OLMo (Groeneveld et al., 2024). In fig. 5 we can see the development of activation outliers early on in the training of both models, although the outliers in OLMo take longer to develop. Furthermore, we confirm a pattern found across the literature, that the primary place where outliers develop is not between layers in a given attention or MLP block but in the residual stream between blocks. That is, the types of layers that do or do not develop outliers are the same in both our model and the pretrained models (e.g., QKV Input activations have outlier channels, while MLP Proj Input activations do not).
4 Mitigating Outlier Channels with Activation Regularization
Based on insights from the previous section, we propose a simple regularization strategy for quantizing the activations of the linear layers, where we use QAT on the input activations and simultaneously penalize the kurtosis of the layer’s outputs.
4.1 Input Activations: QAT with Learned Clip Values
As evident from §2.1, the clip values $c^{-}$ and $c^{+}$ play a key role in uniform quantization. Following PACT (Choi et al., 2018) and LSQ (Bhalgat et al., 2020), we treat these quantization parameters as learnable parameters and optimize them with gradient descent. Concretely, during the forward pass we run the quantization/dequantization step, as shown in algorithm 1. For the backward pass, we use a straight-through estimator to obtain $\nabla{\mathbf{A}}$, $\nabla c^{+}$, $\nabla c^{-}$ from $\nabla\widehat{\mathbf{A}}$ (the gradients with respect to the quantized/dequantized layer). This is shown in algorithm 2. We will show in our experiments that quantizing during training is crucial for 4-bit quantization; just clamping the activations without quantization leads to poor performance.
4.2 Output Activations: Kurtosis Regularization
In our initial experiments we found that QAT on a layer’s input is sufficient to train a W16A4 model that matches the performance of a W16A16. However, since we do not perform QAT for the weights, efficient deployment requires post-training weight quantization to 4 bits. While existing work has shown that weight-only PTQ to 4 bits (i.e., W16A16 $\to$ W4A16) can be done almost losslessly (Frantar et al., 2022; Shao et al., 2023), we observed this to not be the case with QAT models, with W16A4 $\to$ W4A4 resulting in nontrivial perplexity degradations. This is due to the fact that a model can essentially “migrate” the outlier channels to the corresponding rows of the weight matrix, which makes per-column weight PTQ more difficult (as shown in fig. 1(b), bottom).
Algorithm 1 QAT forward pass
0:  $\mathbf{A}$, $c^{-}$, $c^{+}$, $b$, align_zero;
  $s=\frac{2^{b}-1}{c^{+}-c^{-}}$
  if align_zero then
     $z=\operatorname{round}(s\times c^{-})$
  else
     $z=0$
  end if
  $\mathbf{Q}_{\mathbf{A}}=\operatorname{round}(s\times\text{ clamp }(\mathbf{A},c^{-},c^{+})+z)$
  $\widehat{\mathbf{A}}=\frac{1}{s}(\mathbf{Q}_{\mathbf{A}}-z)$
  return  $\widehat{\mathbf{A}}$
Figure 6: The forward and backward passes of QAT. Here $\mathbf{A}$ is the activation tensor, $b$ is the bit width, $c^{-}$ and $c^{+}$ are the learned clip values, and $\nabla\widehat{\mathbf{A}}$ is the gradient with respect to $\widehat{\mathbf{A}}$.
Algorithm 2 QAT backward pass
0:  $\mathbf{A}$, $c^{-}$, $c^{+}$, $b$, $s$, $\nabla\widehat{\mathbf{A}}$;
  $\mathbf{Q}=s\times(\mathbf{A}-c^{-})$
  $\mathbf{E}=(\mathbf{{Q}}-\operatorname{round}(\mathbf{{Q}}))/(2^{b}-1)$
  $\nabla\mathbf{A}_{ij}=\begin{cases}0&\text{if }\mathbf{A}_{ij}>c^{+}\text{ or }\mathbf{A}_{ij}<c^{-}\\ \nabla\widehat{\mathbf{A}}_{ij}&\text{otherwise}\end{cases}$
  $\mathbf{C}^{+}_{ij}=\begin{cases}\nabla\widehat{\mathbf{A}}_{ij}&\text{if }\mathbf{A}_{ij}>c^{+}\\ -\ \mathbf{E}_{ij}\times\nabla\widehat{\mathbf{A}}_{ij}&\text{elif }\mathbf{A}_{ij}>c^{-}\\ 0&\text{otherwise}\end{cases}$
  $\mathbf{C}^{-}_{ij}=\begin{cases}-\ \mathbf{E}_{ij}\times\nabla\widehat{\mathbf{A}}_{ij}&\text{if }\mathbf{A}_{ij}<c^{+}\\ \nabla\widehat{\mathbf{A}}_{ij}&\text{elif }\mathbf{A}_{ij}<c^{-}\\ 0&\text{otherwise}\end{cases}$
  $\nabla c^{+}=\sum_{ij}\mathbf{C}^{+}_{ij}$,       $\nabla c^{-}=\sum_{ij}\mathbf{C}^{-}_{ij}$
  return  $\nabla\mathbf{A},\nabla c^{+},\nabla c^{-}$
One approach to mitigating these outlier weights would be to directly regularize the weights via QAT or some other approach (e.g., $\ell_{\infty}$-norm regularization). However, we found these direct regularization approaches to result in much worse performance and/or unstable training. We thus adopt a more indirect regularization strategy, exploiting the fact that high input channel weights typically lead to a layer’s outputs having outliers, i.e., the output distribution is heavy-tailed (see fig. 1). Our approach thus regularizes the output distribution’s kurtosis. which measures how heavy-tailed a distribution is. An estimate of the kurtosis of a set of values $\mathbf{x}\in\mathbb{R}^{d}$ is given by, $\operatorname{Kurtosis}(\mathbf{d})=\frac{\sum_{i}^{k}(\mathbf{x}_{i}-\mu)^{4}}{\sigma^{4}+\epsilon},$ where $\mu$ and $\sigma$ are respectively the empirical mean and standard deviation of $\mathbf{x}$, and $\epsilon$ is a small term for numerical stability. We multiply the sum of the kurtosis estimates for each token with hyperparameter $\lambda$, and add the result to the cross-entropy loss. While prior work has shown the benefits of regularizing the kurtosis of a layer’s activation distribution to be close to that of a uniform distribution (Chmiel et al., 2020), regularizing the output distribution’s kurtosis to make it less heavy-tailed has not been explored before to our knowledge.
4.3 Post-training Weight Quantization
After training the model to W16A4 with activation regularization on both the inputs/outputs, we experiment with two methods for quantizating the weights to 4 bits. The simplest baseline we use is round-to-nearest (RTN) quantization, which for our purposes implies per-token (for activations) [^4] While there are more sophisticated activiation quantization approaches (Yuan et al., 2023; Chee et al., 2023), these typically have additional overhead (for low-precision matmuls) and are thus not as fast as simple RTN integer quantization. or per-output-channel (for weights) uniform min-max quantization. While the underperformance of RTN weight quantization versus more sophisticated quantization strategies that use calibration data is widely known, we deliberately include this simple data-agnostic baseline to show that activation regularization results in weights that are also easier to quantize (i.e., less perplexity degradation with RTN). Our second approach applies GPTQ (Frantar et al., 2022), which uses a small amount of calibration data to quantize the weights, and is still near the state-of-the-art for 4-bit weight quantization.
5 Empirical Study
Native Activations 4-bit Activations
Weight Precision 16 4 4 4 3 3
Weight Quantizer None GPTQ GPTQ RTN GPTQ RTN
Baseline 23.57 24.10 113233 11855 11755 17187
Activation Clamping 23.73 24.85 378 423 568 663
Kurtosis Regularization 23.72 24.57 8720 8140 10235 19665
QAT 24.30 25.32 25.32 27.76 32.56 46.47
QAT + Kurtosis Regularization 24.10 24.57 24.57 24.90 26.83 30.46
Baseline 25.70 26.16 8430 10028 9107 14498
Activation Clamping 26.38 27.60 32378 6852 26120 15908
Kurtosis Regularization 26.28 26.95 7319 6852 9066 15908
QAT 26.72 27.86 27.87 32.70 64.61 58.81
QAT + Kurtosis Regularization 26.11 26.56 26.56 27.13 30.12 33.46
Table 1: Perplexity of 1B models on C4 (top) and PTB (bottom). Native activation are 16 bits for Baseline, Activation Clamping, Kurtosis Regularization; and 4 bits for QAT, QAT + Kurtosis Regularization.
5.1 Experimental Setup
We use the Megatron-LM (Shoeybi et al., 2020) codebase and train on the SlimPajama dataset (Soboleva et al., 2023). While the trajectory analyses in §3 were done for 50B tokens, due to limited compute we train for 20B tokens for these experiments.
Baselines.
In order to isolate the contributions of each component of our method, we compare against several baselines, on top of the standard-precision baseline. The activation clamping baseline uses static, per-layer clipping values to clamp the input activations. To advantage this approach as much as possible, we use an “oracle” clipping values obtained from QAT to decide the per-layer clipping values, which was found to be more effective than grid-searching on the clipping values. In activation clamping the activations are not quantized during training, and thus this baseline isolates the effect of QAT. The kurtosis regularization baseline applies kurtosis regularization just on the outputs, without QAT. The QAT-only baseline just applies QAT in the input activations.
Hyperparameters.
All hyperparameters were tuned for our 1B W16A16 baseline and kept constant throughout experiments, except for weight decay where we selected between $\{0.1,0.01\}$ for all methods. We use a batch size of 1M tokens, learning rate of 1.5e-4, cosine learning decay, and FP16 precision. For QAT we initialize our clipping values to ±$\pm 4$ for clipping value initializations, unless the layer’s input is bounded. We use the same learning rate but no momentum or weight decay for clip values. For kurtosis we use 1e-5 as the regularization strength.
Evaluation.
We evaluate the perplexity of each model on the C4 and PTB datasets. We test models in three different weight quantization categories: 16 bits, 4 bits, and 3 bits. The 4-bit and 3-bit experiments test with both RTN and GPTQ. For activations, we test in native precision (16 bits for non-QAT models, and 4 bits for the QAT models) as well as in 4 bits. For GPTQ we use a small amount of C4 data for calibration.
5.2 Results
We report the results of our 1B experiments on the C4 and PTB dataset in table 1. We observe that our approach can learn a W4A4 model that has respectable performance compared to the W16A16 baseline. We also observe that the gap between the QAT model with and without kurtosis expands as weights are quantized more and more. At full precision, the gap is less than 1%. At 4 bits, this expands to between 3% and 4%, and at 3 bits this gap widens to 21%. All non-QAT method have catastrophic performance degradations with 4-bit activations. Activation clamping is the only method that achieves less than two orders of magnitude increase in perplexity. In table 2 we perform experiments on downstream tasks for select models to validate our usage of perplexity as a proxy for downstream performance. We observe that models with similar perplexity exhibit similar downstream performance.
We also perform a suite of experiments at the 300M scale, where we just experiment with the QAT baselines. This is shown in  table 3. We largely observe the same trends, with one exception: the gap between the QAT and QAT+Kurtosis Regularization model is smaller than at the 1B scale.
Model Setting HellaSwag PIQA ARC-easy
Baseline W16A16 32.13% 65.51% 48.32%
QAT W16A4 31.79% 65.56% 47.85%
QAT + Kurtosis Regularization W16A4 31.50% 64.96% 48.36%
Table 2: Downstream evaluation of our 1B models on HellaSwag, PIQA, and ARC-easy.
Native Activations 4b Activations
Weight Precision 16 4 4 4 3 3
Weight Quantizer None GPTQ GPTQ RTN GPTQ RTN
Baseline 29.23 30.36 4288 3864 4820.5 3923.96
QAT 30.25 31.30 31.30 32.55 36.47 44.73
QAT+Kurtosis Regularization 29.95 30.83 30.83 31.73 35.47 45.04
Baseline 32.61 34.12 2974 2896 3767 2950
QAT 33.56 34.83 34.83 34.24 47.51 51.22
QAT+ Kurtosis Regularization 33.14 34.23 34.23 34.55 40.74 52.63
Table 3: Perplexity of 300M models on C4 (top) and PTB (bottom). Native activation are 16 bits for Baseline, Activation Clamping, Kurtosis Regularization; and 4 bits for QAT, QAT + Kurtosis Regularization.
5.3 Analysis
Post-Training Quantization of Activations.
Our method shows that QAT from scratch is effective for training a model with 4-bit activations. However, given that most available pretrained models are not trained with 4-bit activations, it would be ideal if we could take a 16-bit activation model and then finetune it with QAT to 4 bits. To test for whether this is possible, we performed an extensive hyperparameter search for QAT finetuning on the pretrained 300M baseline model, where we finetune with QAT for 1B tokens. Even with extensive hyperparameter tuning, QAT finetuning resulted in a W4A4 model with a 16% degradation in perplexity over the W16A16 baseline. Upon further investigation, we found that while our QAT-pretrained models were able to learn to clip outliers without hurting performance, the QAT finetuning models struggled to do so. Finetuning the model longer than 1 billion tokens did not improve results.
We also tried applying OmniQuant (Shao et al., 2023), a state-of-the-art weight-and-activation method for PTQ, to go from W16A16 to W4A4. We found this approach to not perform well, with a significant degradation in perplexity with the 1B model (74.99 on C4 and 107.29 on PTB). Our degradation is larger than what has been reported for pretrained models in the original paper, which could potentially be due to our use of a smaller model (which are typically harder to quantize). Given that the outlier channels seem to emerge early in training (§3), these negative results highlight the importance of early-training interventions for achieving 4-bit activation models.
Direct Approaches for Weight Regularization.
Our use of kurtosis regularization on the output activations to mitigate the effect of “quantization difficulty migration” from the activations to the weights is admittedly indirect. We also experimented with more direct methods for controlling the outliers in the weights: regularizing the kurtosis of the weights instead (at the tensor-level or at the column-level); and regularizing the weight’s $l_{\infty}$ norm. Despite an extensive hyperparameter search, these methods led to unstable training, and we were unable to get these models to converge (unless the regularization-strength hyperparameter was so low that there was effectively no regularization). QAT on the weights also proved unsuccesful, with QAT-weight models underperforming baselines by a significant margin.
Throughput.
Our QAT approach requires modifying the forward and backward passes, which adds nontrivial overhead with an unoptimized, torch.compile-only implementation. This is mainly due to the reduction step in the clip val gradient in the backward pass. We thus implemented our own CUDA kernels that perform a blockwise reduction followed by atomic additions to enable faster throughput. The throughput of our custom kernels on a single H100 node (with eight GPUs) is shown in table 4. We find that while there is still some reduction in throughput, it is closer to the baseline setting than the torch.compile implementation. Given that the numbers in table 4 are from a single node, we anticipate that the actual throughput differences would be even smaller when taking into account the necessary overheads of distributed training.
6 Limitations & Discussion
There are several limitations to our study. While we experiment with language modeling at moderate scale, we were unable to perform experiments on larger models (and train for longer) due to limited compute resources. However, we note that while the 300M parameter models did not benefit as much from the kurtosis intervention on top of QAT, at 1B there was quite a large benefit; this gives us optimism for the utility of our methods at larger scale.
Our study targets integer quantization to 4 bits to enable the use of INT4 matmuls, which is supported by the Ampere architecture GPUs. The more recent GPU architectures (Hopper, Blackwell) unfortunately do not natively support INT4 matmuls, which limit the applicability of our approach on these GPUs. However, the latest Blackwell architecture supports FP4 computations, [^5] https://www.nvidia.com/en-us/data-center/technologies/blackwell-architecture/ and it is possible that QAT may improve FP4-training and moreover enable even lower-precision quantization.
Finally, our study focuses on quantizing only the activations of inputs to linear layers, since linear matmuls consumes the majority of FLOPs during LLM inference (on moderate-length sequences). Future work could consider applying QAT to quantize the activations involved in the attention computations, which could be extremely useful in long-context settings.
Model size Batch size Baseline QAT (torch.compile) QAT (our custom CUDA kernel)
1B 1M tokens 41913 20195 37510
3B 2M tokens 15161 7519 13142
Table 4: Throughput in terms of tokens per second (TPS) on a single node with eight H100s (higher is better). The baseline achieves approximately 50% mean FLOPs utilization (MFU), while our kernel achieves 45%.
7 Conclusion
We study outlier channels in language models from a pretraining perspective. We show that these channels emerge early in pretraining, and are moreover particularly numerous in activations with residual streams. Based on these findings, we propose a simple strategy for mitigating the effect of these outlier channels through activation regularization. We regularize the input activations with QAT plus learned clip values, and we further regularize the output activations via the kurtosis. Our approach is able to learn a W4A4 language model at reasonable scale (1 billion parameters trained on 20B tokens) that is competitive with the standard-precision W16A16 baseline.
Acknowledgments
This study was supported by funds from an MIT-IBM Watson AI grant.